OpenStudy (anonymous):

The surface area of a rectangular prism is 150 cm². If all the sides are tripled, what will the new surface area be?

OpenStudy (sumeer):

the area can be written as 2ab+2ac+2bc=150, the the new area will be 2.3.a.3.b+2.3.a.b+2.3b.3.c=9(2ab+2ac+2bc) => 9*150=1350